OpenStudy (anonymous):

Help??? Assuming x >0, what is the simplified fraction form for 15x times sqrt(3) all divided by sqrt(75x^2)?

OpenStudy (anonymous):

\[\frac{ 15x \sqrt{3} }{ \sqrt{ 75 x^2} }\]

OpenStudy (anonymous):

I think that its 3/1 is that it?

OpenStudy (phi):

I would do it step by step. do you know how to simplify sort(x^2) ?

OpenStudy (phi):

*sqrt(x^2)

OpenStudy (phi):

\[ \sqrt{x^2} = x \] or \[ \sqrt{x \cdot x} = x \] if you have a pair multiplied together inside a square root sign, you can replace it with a single copy with no root sign. so \[ \sqrt{2\cdot 2} = 2 \] or \[ \sqrt{a\cdot a} = a \]

OpenStudy (phi):

\( x^2 \) is short-hand for \(x\cdot x\) so what is \(\sqrt{x^2} \)?

OpenStudy (anonymous):

\[\sqrt{x^2}\] is equal to x

OpenStudy (phi):

yes, so the fraction is now \[ \frac{15x\sqrt{3}}{x\sqrt{75} } \] x divided by x is one, so you can "cancel" them \[ \frac{15\cancel{x}\sqrt{3}}{\cancel{x}\sqrt{75} } = \frac{15\sqrt{3}}{\sqrt{75} } \]

OpenStudy (phi):

now to simplify the \(\sqrt{75}\) break up the 75 into its prime factors 75= 3*5*5 \[ 75= \sqrt{3⋅5⋅5}\] do you see any pairs that you can "pull out" of the square root?

OpenStudy (phi):

Just noticed 5*5*5=125 we want 3*5*5= 75

OpenStudy (anonymous):

Sorry I was gone, I to have dinner. Okay so can you take out the pair of 5s ?

OpenStudy (phi):

yes, the bottom becomes \(5 \sqrt{3} \)

OpenStudy (phi):

\[\frac{15\sqrt{3}}{5\sqrt{3}} \]

OpenStudy (anonymous):

do the square roots of 3 cancel out on the numerator and denominator?

OpenStudy (phi):

yes anything divided by itself (except 0/0) is 1 it is a very useful thing to know.

OpenStudy (anonymous):

okay and then 15/5 is equal to 3/1 or just 3 right?

OpenStudy (anonymous):

Awesome! Thank you so much for helping me do this problem, The steps are clearer now and it will make similar problems easier for me!